# CultureGuard: Towards Culturally-Aware Dataset and Guard Model for Multilingual Safety Applications

### v1

@@ Abstract
Raviraj Joshithanks: Corresponding author., Rakesh Paul, Kanishk Singla, Anusha Kamath, Michael Evans, Katherine Luna, Shaona Ghosh, Utkarsh Vaidya, Eileen Long, Sanjay Singh Chauhan, Niranjan Wartikar NVIDIA {ravirajj, rapaul, kanishks, anushak, michaele, kluna shaonag, uvaidya, elong, schauhan, nwartikar}@nvidia.com
@@ 1 Introduction
Early endeavors in this domain primarily concentrated on multilingual safety guard evaluation datasets Deng et al. ; de Wynter et al. (2025). While models like LlamaGuard Chi et al. (2024) and ShieldGemma Zeng et al. (2024) offer some support for non-English languages, their performance is often suboptimal compared to their English counterparts. Furthermore, the proprietary nature of their datasets and methodologies restricts their utility and customizability for broader research and application. More recent contributions, such as PolyGuard Kumar et al. (2025), DuoGuard Deng et al. (2025), and OmniGuard Verma et al. (2025), have emerged. While PolyGuard is a valuable contribution to multilingual safety, its dependence on labels generated by GPT-4o could raise issues regarding its commercial viability. OmniGuard, on the other hand, is a model-specific solution that presents practical challenges for broader applicability. DuoGuard supports only four languages and does not provide consistent performance across benchmarks. Moreover, these models predominantly leverage machine translation without adequately addressing the crucial aspect of cultural relevance, which is paramount for effective cross-cultural safety.
In response to these challenges, we present CultureGuard, a novel framework designed for curating culturally aligned, high-quality safety datasets across multiple languages. This method is purely based on synthetic data generation, leveraging the English generation capabilities of LLMs and strong translation methods. Importantly, it does not require human input, which presents a scalable approach for targeting multiple languages or regions. Our technical approach employs a four-stage pipeline, encompassing cultural data segregation, cultural data adaptation, machine translation, and quality filtering. Cultural adaptation of the samples using LLMs to the target culture is central to the CultureGuard approach; its core components are shown in Figure 1. It is important to note that since this adaptation is performed on the English source text prior to translation, our approach focuses on ideational and social cultural elements, while linguistic aspects are not explicitly covered Liu et al. (2025). We also present a novel cross-lingual consistency filter to remove low-quality samples post-translation. This pipeline facilitates the systematic conversion and expansion of the English Nemotron-Content-Safety-Dataset-V2 (formerly known as Aegis 2.0) safety dataset Ghosh et al. (2025) into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese. Additionally, we synthetically curate jail-breaking prompt-response pairs using LLMs to enhance the detection capabilities for such adversarial inputs. Safety labels for these synthetic pairs are computed via a jury of LLMs. The final dataset, termed Nemotron-Content-Safety-Dataset-Multilingual-v1, consists of a total of 386,661 samples across 9 languages (including English) as shown in Figure 2. The dataset is based on English Nemotron-Content-Safety-Dataset-V2 and uses the same safety risk taxonomy.
This comprehensive dataset is then used to fine-tune the multilingual Llama-3.1-8B-Instruct model using a LoRA-based approach, resulting in the Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 model. The empirical evaluation of this model quantitatively demonstrates its efficacy, outperforming all publicly available LLMs on a diverse multilingual test dataset. Specifically, our fine-tuned model achieves a 30.2% improvement in multilingual scores compared to the publicly released Llama-Nemotron-Safety-Guard-V2 model (formerly known as Llama-3.1-AegisGuard) and approximately 1.3% over PolyGuard-Qwen, despite PolyGuard’s dataset being three times larger and non-commercially usable. This establishes Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 as the sole multilingual safety guard model currently available with a commercially-friendly license. This marks a significant stride towards bridging the safety gap in multilingual LLMs by enabling the development of truly culturally-aware safety guard models.
@@
  • This dataset is used to train Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1, which is a state-of-the-art model outperforming all other multilingual LLMs and Guard models on multiple multilingual benchmarks. The dataset and model will be released publicly.
@@
For clarity and simplicity throughout this paper, we adopt the following terminology. The suite of models trained in this work are collectively referred to as the CultureGuard model variants. Among these, we identify Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 as the best-performing variant, which utilizes all the components proposed in this work. The dataset, Nemotron-Content-Safety-Dataset-Multilingual-v1, will be referred to as the CultureGuard dataset. More broadly, the name CultureGuard also represents our proposed pipeline for curating such a multilingual cultural dataset.
@@ 2 Related Work
Recent progress in multilingual content safety has centered around the development of comprehensive datasets and robust evaluation methodologies. Key benchmarks include XSafety Wang et al. (2024), MultiJail Deng et al. , RTP-LX de Wynter et al. (2025), and the PolygloToxicityPrompts (PTP) dataset Jain et al. (2024). These resources typically combine naturally occurring human–LLM interactions with human-verified machine translations to provide a more holistic assessment of LLM safety capabilities beyond English. They evaluate a range of dimensions, from general harmfulness and toxicity to jailbreak susceptibility and culturally specific safety violations.
@@ 3.1.2 Cultural Data Adaptation
After adaptation, a jury of LLMs assigns safety labels to the culturally adjusted data. Re-evaluating the safety label is critical because the cultural adaptation process might inadvertently alter the safety profile of the text. To ensure the integrity of the safety labels, we employ a diverse set of jury LLMs including Llama-Nemotron-Safety-Guard-V2111Llama-Nemotron-Safety-Guard-V2, PolyGuard-Qwen [^2] PolyGuard-Qwen, Llama-3.1-Nemotron-70B-Instruct [^3] Llama-3.1-Nemotron-70B-Instruct, and Qwen3-235B-A22B [^4] Qwen3-235B-A22B; we use the FP8-quantized version throughout this work. (non-reasoning). An adapted sample is retained only if its majority-voted safety label aligns with the ground truth label of the original text.
@@ 3.1.4 Quality Filtering
Models Average CultureGuard CultureGuard PGPrompts PGPrompts RTP-LX MultiJail XSafety (Prompts) (Response) (Prompts) (Response) Nemotron-Safety-Guard-V2 63.26 63.92 78.96 57.62 68.50 66.45 72.08 35.31 Llama-Guard-3-8B 61.46 71.78 60.54 67.34 64.68 47.01 78.04 40.82 Llama-Guard-4-12B 56.02 67.47 57.79 62.02 53.08 39.12 73.56 39.13 PolyGuard-Qwen 81.34 84.94 80.41 84.68 74.19 86.98 93.67 64.50 CultureGuard (vanilla) 79.96 84.84 85.46 74.85 71.38 90.39 92.91 59.88 CultureGuard 81.04 85.58 85.56 78.88 70.82 89.87 93.69 62.90 CultureGuard + JB SDG 81.35 85.43 84.57 79.35 73.44 89.97 93.94 62.77 CultureGuard + JB SDG + QF 82.37 85.15 85.48 79.25 72.89 91.49 95.36 66.97
@@
Models CultureGuard-JB CultureGuard-JB (Prompt) (Response) Nemotron-Safety-Guard-V2 66.73 86.55 Llama-Guard-3-8B 72.29 77.66 Llama-Guard-4-12B 69.63 70.14 PolyGuard-Qwen 88.97 89.89 CultureGuard (vanilla) 86.78 91.23 CultureGuard 88.53 91.99 CultureGuard + JB SDG 91.54 94.64 CultureGuard + JB SDG + QF 91.77 94.35
@@ 4 Results and Discussion
This section presents the performance evaluation of various safety guard models across a diverse set of multilingual benchmarks. The harmful-f1 score is the metric used to compare the models. Table 1 shows the summary of results averaged across 9 languages. For full results for all languages, refer to the Appendix A. We evaluate public Guard models including Llama-Nemotron-Safety-Guard-V2, Llama-Guard-3-8B, Llama-Guard-4-12B, and PolyGuard-Qwen, in comparison to our proposed CultureGuard variants on the CultureGuard dataset, PolyGuardPrompts (PGPrompts), RTP-LX, MultiJail, and XSafety datasets. Our findings demonstrate the superior efficacy of CultureGuard, particularly when augmented with Jail-Breaking Synthetic Data Generation (JB SDG) and Quality Filtering (QF). We also present ablations to show the impact of adding synthetic jail-breaking data and the quality filtering step. The following CultureGuard variants, also referred to in Table 1, represent different ablations:
@@
Models CultureGuard CultureGuard (Prompts) (Response) Guard Models CultureGuard + JB SDG + QF 85.15 85.48 Nemotron-Safety-Guard-V2 63.92 78.96 PolyGuard-Qwen 84.94 80.41 Llama-Guard-3-8B 71.78 60.54 Llama-Guard-4-12B 67.47 57.79 Generic SLMs Llama-3.2-3B-Instruct 78.14 71.29 Llama-3.1-Nemotron-Nano-4B-v1.1 (reasoning off) 68.38 68.27 Llama-3.1-Nemotron-Nano-4B-v1.1 (reasoning on) 72.35 66.61 Llama-3.1-8B-Instruct 77.56 78.54 Qwen3-8B (reasoning off) 77.22 81.81 Qwen3-8B (reasoning on) 79.60 75.71 Generic LLMs Gemma-2-27b-it 82.82 83.03 Gemma-3-27b-it 82.03 79.29 Qwen3-32B (reasoning off) 82.78 81.12 Qwen3-32B (reasoning on) 83.63 80.01 Llama-3.3-Nemotron-Super-49B-v1 (reasoning off) 81.33 79.16 Llama-3.3-Nemotron-Super-49B-v1 (reasoning on) 83.24 76.51 Llama-3.3-70B-Instruct 83.55 82.14 Qwen3-235B-A22B (reasoning off) 82.95 81.33 Qwen3-235B-A22B (reasoning on) 83.98 78.69 Llama-3.1-405B [^5] Model refusals to categorize samples due to safety concerns were treated as unsafe classifications. 84.60 83.56
Table 3: Comparison of model performance on CultureGuard test-set prompts and responses (average across languages). The generic SLMs and LLMs [^7] Closed models like GPT-4o are omitted as their API safety filters interfered with the evaluation, blocking the required test prompts. act as Guard models using the same CultureGuard prompt during inference.
@@ 4.1 Impact of Multilingual Data
The English Llama-Nemotron-Safety-Guard-V2 model, despite having the same backbone as its multilingual counterpart, generally exhibits lower performance compared to CultureGuard (vanilla) and the CultureGuard model. This highlights the importance of incorporating multilingual data for robust safety performance across languages. For instance, the CultureGuard (vanilla) model achieves an average performance of 79.96, significantly higher than English Llama-Nemotron-Safety-Guard-V2’s 63.26.
@@ 4.2 Impact of Cultural Adaptation
CultureGuard model consistently demonstrates superior performance compared to the CultureGuard (vanilla) model. For instance, the CultureGuard model achieves an average performance of 81.04, while the CultureGuard (vanilla) model scores 79.96. This shows the importance of cultural adaptation in our approach, as the CultureGuard (vanilla) model utilizes a vanilla translation and filtering recipe without explicit cultural alignment. On the CultureGuard test set, the CultureGuard model (without JB SDG and QF) demonstrates strong performance, achieving the highest scores for both prompts and responses. PolyGuard-Qwen also shows competitive results on these internal benchmarks.
@@ 4.3 Impact of Jail-Breaking Synthetic Data and Filtering
The CultureGuard + JB SDG + QF model variant achieves the highest overall average performance of 82.37, indicating its robust and generalized capability in multilingual safety detection.
For external benchmarks, CultureGuard + JB SDG + QF model consistently leads, achieving the highest scores on RTP-LX, MultiJail, and XSafety. This shows the effectiveness of incorporating jail-breaking synthetic data and quality filtering. PolyGuard-Qwen, while strong on its native PolyGuardPrompts benchmark, generally demonstrates lower performance than the final CultureGuard variant on other external evaluations. The incremental improvements across CultureGuard variants underscore the positive impact of our methodology stages. On the synthetically curated CultureGuard-JB test set (jail-breaking prompts) as well, our model performs better than other competitive models, showcasing its superior ability to classify jail-breaking prompts (Table 2).

### v2

@@ Abstract
Raviraj Joshithanks: Corresponding author., Rakesh Paul, Kanishk Singla, Anusha Kamath, Michael Evans, Katherine Luna, Shaona Ghosh, Utkarsh Vaidya, Eileen Long, Sanjay Singh Chauhan, Niranjan Wartikar NVIDIA {ravirajj, rapaul, kanishks, anushak, michaele, kluna, shaonag, uvaidya, elong, schauhan, nwartikar}@nvidia.com
@@ 1 Introduction
Figure 1: A comparison of a multilingual safety guard model’s performance across different languages.
@@
Early endeavors in this domain primarily concentrated on multilingual safety guard evaluation datasets Deng et al. ; de Wynter et al. (2025). While models like LlamaGuard Chi et al. (2024), ShieldGemma Zeng et al. (2024), and Granite Guardian Padhi et al. (2024) offer some support for non-English languages, their performance is often suboptimal compared to their English counterparts. Furthermore, the proprietary nature of their datasets and methodologies restricts their utility and customizability for broader research and application. Similarly, open models like Llama-Nemotron-Safety-Guard-Defensive-V1 Ghosh et al. (2024), Llama-Nemotron-Safety-Guard-V2 Ghosh et al. (2025), and WildGuard Han et al. (2024) also lack dedicated multilingual support. More recent contributions, such as PolyGuard Kumar et al. (2025), DuoGuard Deng et al. (2025), and OmniGuard Verma et al. (2025), have emerged. While PolyGuard is a valuable contribution to multilingual safety, its dependence on labels generated by GPT-4o could raise issues regarding its commercial viability. OmniGuard, on the other hand, is a model-specific solution that presents practical challenges for broader applicability. DuoGuard supports only four languages and does not provide consistent performance across benchmarks. Moreover, these models predominantly leverage machine translation without adequately addressing the crucial aspect of cultural relevance, which is paramount for effective cross-cultural safety.
@@
In response to these challenges, we present CultureGuard, a novel framework designed for curating culturally aligned, high-quality safety datasets across multiple languages. This method is purely based on synthetic data generation, leveraging the English generation capabilities of LLMs and strong translation methods. Importantly, it does not require human input, which presents a scalable approach for targeting multiple languages or regions. Our technical approach employs a four-stage pipeline, encompassing cultural data segregation, cultural data adaptation, machine translation, and quality filtering. Cultural adaptation of the samples using LLMs to the target culture is central to the CultureGuard approach; its core components are shown in Figure 2. It is important to note that since this adaptation is performed on the English source text prior to translation, our approach focuses on ideational and social cultural elements, while linguistic aspects are not explicitly covered Liu et al. (2025). We also present a novel cross-lingual consistency filter to remove low-quality samples post-translation. This pipeline facilitates the systematic conversion and expansion of the English Nemotron-Content-Safety-Dataset-V2 (formerly known as Aegis 2.0) safety dataset Ghosh et al. (2025) into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese (Mandarin). Additionally, we synthetically curate jail-breaking prompt-response pairs using LLMs to enhance the detection capabilities for such adversarial inputs. Safety labels for these synthetic pairs are computed via a jury of LLMs. The final dataset, termed Nemotron-Content-Safety-Dataset-Multilingual-v1111Nemotron-Content-Safety-Dataset-Multilingual-v1, consists of a total of 386,661 samples across 9 languages (including English) as shown in Figure 3. The dataset is based on the English Nemotron-Content-Safety-Dataset-V2 and uses the same safety risk taxonomy.
This comprehensive dataset is then used to fine-tune the multilingual Llama-3.1-8B-Instruct model using a LoRA-based approach, resulting in the Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1222Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 model. The empirical evaluation of this model quantitatively demonstrates its efficacy, outperforming all publicly available LLMs on a diverse multilingual test dataset. Specifically, our fine-tuned model achieves a 31.06% relative improvement in multilingual scores compared to the publicly released Llama-Nemotron-Safety-Guard-V2 model (formerly known as Llama-3.1-AegisGuard) and approximately 1.28% over PolyGuard-Qwen, despite PolyGuard’s dataset being three times larger and non-commercially usable. Figure 1 illustrates this performance comparison on a per-language basis. This establishes Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 as the sole multilingual safety guard model currently available with a commercially-friendly license. This marks a significant stride towards bridging the safety gap in multilingual LLMs by enabling the development of truly culturally-aware safety guard models.
@@
  • This dataset is used to train Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1, which is a state-of-the-art model outperforming all other multilingual LLMs and Guard models on multiple multilingual benchmarks. The dataset and model are released publicly.
@@
For clarity and simplicity throughout this paper, we adopt the following terminology. The suite of models trained in this work is collectively referred to as the CultureGuard model variants. Among these, we identify Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 as the best-performing variant, which utilizes all the components proposed in this work. The dataset, Nemotron-Content-Safety-Dataset-Multilingual-v1, is referred to as the CultureGuard dataset. More broadly, the name CultureGuard also represents our proposed pipeline for curating such a multilingual cultural dataset.
@@ 2 Related Work
Recent progress in multilingual content safety has centered around the development of comprehensive datasets and robust evaluation methodologies. Key benchmarks include XSafety Wang et al. (2024), MultiJail Deng et al. , RTP-LX de Wynter et al. (2025), PolygloToxicityPrompts (PTP) Jain et al. (2024) and the Aya Red-teaming Ahmadian et al. (2024) datasets. These resources typically combine naturally occurring human–LLM interactions with human-verified machine translations to provide a more holistic assessment of LLM safety capabilities beyond English. They evaluate a range of dimensions, from general harmfulness and toxicity to jailbreak susceptibility and culturally specific safety violations.
@@ 3.1.2 Cultural Data Adaptation
After adaptation, a jury of LLMs assigns safety labels to the culturally adjusted data. Re-evaluating the safety label is critical because the cultural adaptation process might inadvertently alter the safety profile of the text. To ensure the integrity of the safety labels, we employ a diverse set of jury LLMs including Llama-Nemotron-Safety-Guard-V2333Llama-Nemotron-Safety-Guard-V2, PolyGuard-Qwen [^4] PolyGuard-Qwen, Llama-3.1-Nemotron-70B-Instruct [^5] Llama-3.1-Nemotron-70B-Instruct, and Qwen3-235B-A22B [^6] Qwen3-235B-A22B; we use the FP8-quantized version throughout this work. (non-reasoning). An adapted sample is retained only if its majority-voted safety label aligns with the ground truth label of the original text.
@@ 3.1.4 Quality Filtering
Models Average CultureGuard CultureGuard PGPrompts PGPrompts RTP-LX MultiJail XSafety Aya Red-teaming (Prompts) (Response) (Prompts) (Response) (Prompts) (Prompts) (Prompts) (Prompts) Nemotron-Safety-Guard-V2 64.22 63.92 78.96 57.62 68.50 66.45 72.08 35.31 70.95 Llama-Guard-3-8B 63.89 71.78 60.54 67.34 64.68 47.01 78.04 40.82 80.89 Llama-Guard-4-12B 57.77 67.47 57.79 62.02 53.08 39.12 73.56 39.13 70.01 Granite Guardian 3.1 8B 78.66 78.63 75.16 76.87 66.37 86.26 87.21 64.98 93.79 Granite Guardian 3.3 8B (reasoning off) 74.84 75.43 75.86 72.30 67.92 78.23 80.99 56.93 91.08 Granite Guardian 3.3 8B (reasoning on) 74.29 76.87 72.27 68.47 61.22 78.72 85.71 58.91 92.17 PolyGuard-Qwen 83.10 84.94 80.41 84.68 74.19 86.98 93.67 64.50 95.46 CultureGuard (vanilla) 81.69 84.84 85.46 74.85 71.38 90.39 92.91 59.88 93.81 CultureGuard 82.78 85.58 85.56 78.88 70.82 89.87 93.69 62.90 94.90 CultureGuard + JB SDG 83.12 85.43 84.57 79.35 73.44 89.97 93.94 62.77 95.52 CultureGuard + JB SDG + QF 84.17 85.15 85.48 79.25 72.89 91.49 95.36 66.97 96.79
@@
Models CultureGuard-JB CultureGuard-JB (Prompt) (Response) Nemotron-Safety-Guard-V2 66.73 86.55 Llama-Guard-3-8B 72.29 77.66 Llama-Guard-4-12B 69.63 70.14 Granite Guardian 3.1 8B 83.59 80.93 Granite Guardian 3.3 8B (reasoning off) 82.40 83.97 Granite Guardian 3.3 8B (reasoning on) 81.37 78.55 PolyGuard-Qwen 88.97 89.89 CultureGuard (vanilla) 86.78 91.23 CultureGuard 88.53 91.99 CultureGuard + JB SDG 91.54 94.64 CultureGuard + JB SDG + QF 91.77 94.35
@@ 3.3 Training and Evaluation
  • Aya Red-teaming: The Aya red-teaming dataset is a human-annotated collection of harmful prompts designed to identify and categorize harmful content across nine different harm categories. The dataset’s overlapping languages include Arabic, English, French, Hindi, and Spanish.
@@ 4 Results and Discussion
This section presents the performance evaluation of various safety guard models across a diverse set of multilingual benchmarks. The harmful-f1 score is the metric used to compare the models. Table 1 shows the summary of results averaged across 9 languages. For full results for all languages, refer to the Appendix A. We evaluate public Guard models, including Llama-Nemotron-Safety-Guard-V2, Llama-Guard-3-8B, Llama-Guard-4-12B, Granite Guardian 3.1 8B, and PolyGuard-Qwen, in comparison to our proposed CultureGuard variants on the CultureGuard dataset, PolyGuardPrompts (PGPrompts), RTP-LX, MultiJail, XSafety, and Aya Red-teaming datasets. Our findings demonstrate the superior efficacy of CultureGuard, particularly when augmented with Jail-Breaking Synthetic Data Generation (JB SDG) and Quality Filtering (QF). We also present ablations to show the impact of adding synthetic jail-breaking data and the quality filtering step. The following CultureGuard variants, also referred to in Table 1, represent different ablations:
@@
Models CultureGuard CultureGuard (Prompts) (Response) Guard Models CultureGuard + JB SDG + QF 85.15 85.48 Nemotron-Safety-Guard-V2 63.92 78.96 PolyGuard-Qwen 84.94 80.41 Llama-Guard-3-8B 71.78 60.54 Llama-Guard-4-12B 67.47 57.79 Granite Guardian 3.1 8B 78.63 75.16 Granite Guardian 3.3 8B (reasoning off) 75.43 75.86 Granite Guardian 3.3 8B (reasoning on) 76.87 72.27 Generic SLMs Llama-3.2-3B-Instruct 78.14 71.29 Llama-3.1-Nemotron-Nano-4B-v1.1 (reasoning off) 68.38 68.27 Llama-3.1-Nemotron-Nano-4B-v1.1 (reasoning on) 72.35 66.61 Llama-3.1-8B-Instruct 77.56 78.54 Qwen3-8B (reasoning off) 77.22 81.81 Qwen3-8B (reasoning on) 79.60 75.71 Generic LLMs Gemma-2-27b-it 82.82 83.03 Gemma-3-27b-it 82.03 79.29 Qwen3-32B (reasoning off) 82.78 81.12 Qwen3-32B (reasoning on) 83.63 80.01 Llama-3.3-Nemotron-Super-49B-v1 (reasoning off) 81.33 79.16 Llama-3.3-Nemotron-Super-49B-v1 (reasoning on) 83.24 76.51 Llama-3.3-70B-Instruct 83.55 82.14 Qwen3-235B-A22B (reasoning off) 82.95 81.33 Qwen3-235B-A22B (reasoning on) 83.98 78.69 Llama-3.1-405B [^7] Model refusals to categorize samples due to safety concerns were treated as unsafe classifications. 84.60 83.56
Table 3: Comparison of model performance on CultureGuard test-set prompts and responses (average harmful-F1 score across languages). The generic SLMs and LLMs [^9] Closed models like GPT-4o are omitted as their API safety filters interfered with the evaluation, blocking the required test prompts. act as Guard models using the same CultureGuard prompt during inference.
@@ 4.1 Impact of Multilingual Data
The English Llama-Nemotron-Safety-Guard-V2 model, despite having the same backbone as its multilingual counterpart, generally exhibits lower performance compared to CultureGuard (vanilla) and the CultureGuard model. This highlights the importance of incorporating multilingual data for robust safety performance across languages. For instance, the CultureGuard (vanilla) model achieves an average performance of 81.69, significantly higher than English Llama-Nemotron-Safety-Guard-V2’s 64.22.
@@ 4.2 Impact of Cultural Adaptation
CultureGuard model consistently demonstrates superior performance compared to the CultureGuard (vanilla) model. For instance, the CultureGuard model achieves an average performance of 82.78, while the CultureGuard (vanilla) model scores 81.69. This shows the importance of cultural adaptation in our approach, as the CultureGuard (vanilla) model utilizes a vanilla translation and filtering recipe without explicit cultural alignment. On the CultureGuard test set, the CultureGuard model (without JB SDG and QF) demonstrates strong performance, achieving the highest scores for both prompts and responses. PolyGuard-Qwen also shows competitive results on these internal benchmarks.
@@ 4.3 Impact of Jail-Breaking Synthetic Data and Filtering
The CultureGuard + JB SDG + QF model variant achieves the highest overall average performance of 84.17, indicating its robust and generalized capability in multilingual safety detection.
For external benchmarks, CultureGuard + JB SDG + QF model consistently leads, achieving the highest scores on RTP-LX, MultiJail, XSafety, and Aya Red-teaming. This shows the effectiveness of incorporating jail-breaking synthetic data and quality filtering. PolyGuard-Qwen, while strong on its native PolyGuardPrompts benchmark, generally demonstrates lower performance than the final CultureGuard variant on other external evaluations. The incremental improvements across CultureGuard variants underscore the positive impact of our methodology stages. On the synthetically curated CultureGuard-JB test set (jail-breaking prompts) as well, our model performs better than other competitive models, showcasing its superior ability to classify jail-breaking prompts (Table 2).